# PROM: Prioritize Reduction of Multiplications Over Lower Bit-Widths for Efficient CNNs

### v1

@@ 1 Introduction
While computer vision models made remarkable progress in the last decade [16, 20, 39, 41, 7, 43], the increasing computational cost of neural networks raises concerns with regard to their energy consumption and environmental impact, as well as their suitability for deployment on resource-constrained and mobile devices. While recent large-scale Transformer models [7, 43] demonstrate high accuracy in vision tasks, their deployment on smaller devices remains challenging. In contrast, convolutional neural networks (CNNs) continue to be widely used, especially in mobile environments, as they strike a balance between performance, resource requirements and training efficiency.
@@ 4 Evaluating Model Efficiency
We follow prior work [51, 14, 3, 28, 27] and do not include cheap elementwise operations, such as batch normalization, in the energy calculation for a fair comparison across methods.
@@
Hardware Support. We discuss the availability of operations required to run different quantization schemes on commodity hardware. While methods relying on int8 computations are readily supported on modern CPUs and GPUs, lower bit-widths are harder to use without specific accelerator chips. We do not directly compare our method to mixed-precision approaches that learn an independent bit-width for each module in the network [5, 4], as these methods typically do not attempt a hardware-level deployment scenario and do not rely on a fixed-integer format in each layer. In contrast, PROM uses precisely ternary weights for pointwise convolutions and 8-bit weights for depthwise convolutions as well as linear layers, simplifying the hardware instructions required for running inference with a model.
@@ 5 Experiments
In this section, we extensively evaluate our method and provide a comparison to related work based on results on the standard ImageNet ILSVRC 2012 benchmark [38] for image classification. We follow common practice and report our results on the validation set of ImageNet. All of our models are implemented in PyTorch [34] and trained as described in Section 3.2. For the MobileNetV2 architecture [39], we train a suite of models by altering the width multiplier setting, ranging from 0.75 to 2.0. This setting adjusts the width of each module by the given factor. Similarly, we train RegNetX [35] and ResNeXt [50] models with varying depths, allowing us to observe the scaling properties of our method.
@@ 5.1 Results on ImageNet
| Baseline | 0.75MobileNetV2 | 16/16 | 69.15 | 5.27 | 308.7 | 102.6 |
| 1.0MobileNetV2 | 16/16 | 71.78 | 7.01 | 445.4 | 148.1 | |
| 1.25MobileNetV2 | 16/16 | 73.37 | 10.10 | 733.0 | 242.8 | |
| 1.5MobileNetV2 | 16/16 | 74.18 | 13.72 | 1000.4 | 332.9 | |
| 1.75MobileNetV2 | 16/16 | 74.48 | 17.84 | 1326.7 | 441.5 | |
| 2.0MobileNetV2 | 16/16 | 74.66 | 22.52 | 1694.5 | 564.1 | |
| BASQ [24] | 1.0MobileNetV2 | 2/2 | 64.71 | 0.94 | 17.3 | 6.0 |
| PROM | 0.75MobileNetV2 | (1.58/8)/8 | 64.53 | 1.70 | 11.3 | 3.3 |
| PROM | 1.0MobileNetV2 | (1.58/8)/8 | 68.45 | 1.95 | 15.2 | 4.3 |
| GradScale [40] | 1.0MobileNetV2 | 4/4 | 69.77 | 1.80 | 34.4 | 11.7 |
| PROFIT [33] | 1.0MobileNetV2 | 4/4 | 71.56 | 1.80 | 34.4 | 11.7 |
| BASQ [24] | 1.0MobileNetV2 | 4/4 | 71.98 | 1.80 | 34.4 | 11.7 |
| PROM | 1.25MobileNetV2 | (1.58/8)/8 | 71.58 | 2.60 | 24.5 | 6.2 |
| PROM | 1.5MobileNetV2 | (1.58/8)/8 | 73.17 | 3.31 | 29.5 | 8.1 |
| PROFIT [33] | 1.0MobileNetV2 | 8/8 | 72.35 | 3.54 | 68.7 | 23.1 |
| PROM | 1.5MobileNetV2 | (1.58/8)/8 | 73.17 | 3.31 | 29.5 | 8.1 |
| PROM | 1.75MobileNetV2 | (1.58/8)/8 | 74.24 | 4.10 | 37.5 | 10.1 |
| PROM | 2.0MobileNetV2 | (1.58/8)/8 | 74.81 | 4.96 | 46.4 | 12.4 |
@@
| Baseline | RegNetX-400MF | 16/16 | 72.69 | 10.99 | 650.4 | 215.8 |
| RegNetX-800MF | 16/16 | 75.27 | 14.52 | 1308.8 | 434.7 | |
| RegNetX-1.6GF | 16/16 | 76.90 | 18.38 | 2563.9 | 850.8 | |
| LSQ [9] | ResNet18 | 2/2 | 67.60 | 2.94 | 97.2 | 34.0 |
| DSQ [13] | ResNet18 | 4/4 | 69.56 | 10.18 | 193.3 | 66.2 |
| BRECQ [26] | ResNet18 | 4/4 | 69.90 | 5.81 | 193.3 | 66.2 |
| PROM | RegNetX-400MF | (1.58/8)/8 | 68.68 | 2.40 | 33.3 | 10.2 |
| PROFIT [33] | MobileNetV1 | 8/8 | 70.16 | 4.25 | 130.5 | 43.7 |
| LSQ [9] | ResNet18 | 8/8 | 71.10 | 11.69 | 386.6 | 131.1 |
| LSQ [9] | ResNet18 | 4/4 | 71.20 | 5.85 | 193.3 | 66.2 |
| LSQ [9] | ResNet34 | 2/2 | 71.60 | 5.48 | 199.6 | 69.7 |
| PROM | RegNetX-800MF | (1.58/8)/8 | 71.96 | 3.01 | 57.7 | 17.3 |
| DSQ [13] | ResNet34 | 4/4 | 72.76 | 15.24 | 396.8 | 135.8 |
| LSQ [9] | ResNet50 | 2/2 | 73.70 | 6.48 | 278.9 | 96.7 |
| PROM | RegNetX-1.6GF | (1.58/8)/8 | 73.91 | 4.15 | 131.1 | 40.3 |
@@
| Baseline | ResNeXt-50(324d) | 16/16 | 77.42 | 50.04 | 7520.0 | 2504.3 |
| ResNeXt-101(328d) | 16/16 | 79.18 | 177.56 | 25567.5 | 8504.9 | |
| LSQ [9] | ResNet34 | 8/8 | 74.10 | 21.81 | 793.5 | 269.2 |
| LSQ [9] | ResNet34 | 4/4 | 74.10 | 10.92 | 396.8 | 135.8 |
| BRECQ [26] | ResNet50 | 4/4 | 75.05 | 12.85 | 553.6 | 188.4 |
| LSQ [9] | ResNet50 | 4/4 | 76.70 | 12.85 | 553.6 | 188.4 |
| PROM | ResNeXt-50(324d) | (1.58/8)/8 | 76.71 | 8.97 | 195.1 | 51.0 |
| LSQ [9] | ResNet50 | 8/8 | 76.80 | 25.60 | 1107.3 | 372.6 |
| PROM | ResNeXt-101(328d) | (1.58/8)/8 | 78.21 | 32.09 | 878.8 | 249.6 |
By scaling the quantized model’s width (for MobileNetV2) and depth (for RegNetX and ResNeXt), we achieve accurate and highly efficient models for varying storage size and energy budgets. For example, we can reduce the energy consumption of a 1.0MobileNetV2 model by a factor of up to 34.4, and the required memory to store it by 3.6. To nearly match the accuracy of the floating-point 1.0MobileNetV2, our 1.25-scaled model requires 2.7 less storage and 23.9 less energy per forward pass, as depicted in Figure 5. Interestingly, our largest model surpasses the floating-point baseline model’s accuracy by 3.03 percentage points, while using 1.4 less storage and 11.9 less energy on architectures.
We observe that our method scales well with the model’s width or depth. Strong quantization in the form of ternary weights for pointwise convolutions heavily restricts the function approximation properties of the model. This issue can be alleviated by providing it with more weights to tune, increasing its representational power. We also note that the 2.0-scaled MobileNetV2 exhibits a much greater amount of ternary weights equal to zero after training when compared to the 1.0 model. We hypothesize that this indicates an automatic allocation of model capacity during training, essentially "pruning" filters in unimportant layers. As weights which are equal to zero do not need to be computed in the forward pass, the energy expense can be reduced even further on suitable hardware.
@@ 5.2 Energy Evaluation
As visualized in Figure 1, our method improves the Pareto frontier in the trade-off between top-1 accuracy on ImageNet and the energy cost of inference for CNN models. In particular, PROM outperforms other methods for quantizing a MobileNetV2 architecture to lower bit-widths.
As demonstrated in Tables 1, 2 and 3, our method achieves best-in-category energy efficiency on and architectures while offering similar or better performance than prior work. This makes our method well-suited for deployment on battery-powered or resource-constrained devices. Additionally, note that the energy consumption of a model strongly correlates with the ACE metric introduced in [51], which measures a network’s total inference cost on hardware.
@@ 5.4 Comparison to Binary Neural Networks
The concept of low bit-width quantization for convolutional architectures naturally extends to binary networks [36, 21]. Instead of relying on integer addition and multiplication routines, binary networks transform all floating-point operations in a network into logical XNOR and POPCOUNT (counting the number of ones in a bit-string) commands by quantizing both weights and activations into binary tensors.
@@ 5.5 Hardware Support
As quantization is mostly used to decrease the cost of running a model on a given set of hardware in terms of memory and energy requirements, hardware support for the proposed quantization operations is essential. Most modern CPU and GPU architectures include highly optimized instructions for int8 operations such as addition and multiplication, rendering our proposed method efficient and widely applicable across a variety of hardware.
For models with bit-widths of 2-bit [13, 9] and 4-bit [13, 9, 26, 33], an efficient implementation is not straight-forward due to a lack of native hardware support, such as in the popular ARMv8 [2], ARM NEON [1] or modern Intel [22] instruction sets. While specialized hardware could allow these models to run efficiently, general-purpose hardware often lacks optimizations for operations below int8 [49]. This could lead to complicated memory access patterns, as weights need to be packed and unpacked into larger data types, decreasing the model’s efficiency. However, when the hardware support for int4 and int2 operations improves, further reducing the bit-width of the non-ternary components in the network would be an interesting direction for future research.
@@ 6 Related Work
Quantizing Weights and Activations to Integers. The most common approach to quantization is to reduce the bit-width of a model’s weights and activations from 32-bit or 16-bit floating-point values to integers, which includes 8-bit, 4-bit and 2-bit integers. Plenty of related methods [33, 9, 53] explore int8 weights and activations due to the relative ease of use while still yielding good results. Quantizing weights to int4 or int2 precision is more complex and causes larger drops in performance due to reduced model capacity and training instabilities [33, 53, 13]. LSQ [9] proposes a learnable scale factor to the gradients of the task loss, aiming to improve quantization-aware training, especially for low bit-widths. BRECQ [26] employs second-order error analysis to improve post-training quantization. PROFIT [33] freezes specific network layers to mitigate training instabilities. DSQ [13] asymptotically approximates the targeted quantization scheme during a multi-stage training process.
While mixed-precision quantization approaches exist [5, 4], they are not suited to run on general-purpose hardware, as all blocks can have different bit-widths. Furthermore, they are unable to reduce the number of multiplications, as is possible with our proposed method.
Recent work also explores the usage of binary or ternary weights while maintaining high bit-width activations [44, 29, 54]. The forward pass of a layer quantized in this fashion reduces to a simple addition of input channels, completely eliminating multiplications and therefore drastically reducing cost. BitNet [44] and BitNet b1.58 [29] introduce simple and efficient large language models by using linear layers with binary and ternary weights, respectively. This concept is extended to the attention mechanism in [54], eliminating all matrix multiplications from a language model.
@@
Model Compression via Pruning. Aside from quantization, another prominent way to compress models for usage on resource-constrained hardware is pruning [48, 17]. While early magnitude-based schemes [15] eliminate low-importance weights from pre-trained architectures in an unstructured fashion, recent approaches leverage an interspace filter basis [47] or combine different pruning scores [46] to identify and remove the most redundant weights. To achieve cost reduction on general-purpose hardware, structured pruning [45, 25, 12, 8, 32] can be used, removing entire filters from convolution modules or features from latent feature maps. Static approaches [45, 25] train a fixed architecture for every compute budget. In contrast, dynamic approaches [12, 8] learn gating modules, dynamically toggling filters or features depending on the complexity of the input image. Recent approaches even show potential for instant model compression without requiring any training [32].
Pruning reduces the parameter count and inference cost of models, and — when paired with integer quantization — can yield even greater efficiency gains. While identifying structural redundancies within heavily-quantized networks is beyond the scope of this work, it represents an interesting avenue for future research.

### v2

@@ 1 Introduction
While computer vision models have made remarkable progress in the last decade [17, 21, 40, 42], their increasing computational cost raises concerns about energy consumption, environmental impact, and suitability for deployment on resource-constrained devices. This is particularly challenging for large-scale Transformer models [7, 44]. In contrast, convolutional neural networks (CNNs) remain widely used, especially in mobile environments, as they strike a balance between performance, resource requirements and training efficiency.
@@ 4 Evaluating Model Efficiency
We follow prior work [55, 14, 3, 29, 28] and do not include cheap elementwise operations, such as batch normalization [24], in the energy calculation for a fair comparison across methods.
@@
Hardware Support. We discuss the availability of operations required to run different quantization schemes on commodity hardware. While methods relying on int8 computations are readily supported on modern CPUs and GPUs, lower bit-widths are harder to use without specific accelerator chips. Similarly, mixed-precision approaches [5, 4, 54, 47] that learn an independent bit-width for each module in the network complicate hardware deployment, as they do not rely on a fixed quantization format across layers. In contrast, PROM uses precisely ternary weights for pointwise convolutions and 8-bit weights for depthwise convolutions as well as linear layers, simplifying the hardware instructions required for running inference with a model.
@@ 5 Experiments
In this section, we extensively evaluate our method and provide a comparison to related work based on results on the standard ImageNet ILSVRC 2012 benchmark [39] for image classification. We follow common practice and report our results on the validation set of ImageNet. All of our models are implemented in PyTorch [35] and trained as described in Section 3.2. Additionally, we include pseudocode for our implementation in the Supplementary Material (see Figures 10-12). For the MobileNetV2 architecture [40], we train a suite of models by altering the width multiplier setting, ranging from 0.75 to 2.0. This setting adjusts the width of each module by the given factor. Similarly, we train RegNetX [36] and ResNeXt [53] models with varying depths, allowing us to observe the scaling properties of our method.
@@ 5.1 Results on ImageNet
| Baseline | 0.75MobileNetV2 | 16/16 | 69.15 | 5.27 | 308.7 | 102.6 | ||||
| 1.0MobileNetV2 | 16/16 | 71.78 | 7.01 | 445.4 | 148.1 | |||||
| 1.25MobileNetV2 | 16/16 | 73.37 | 10.10 | 733.0 | 242.8 | |||||
| 1.5MobileNetV2 | 16/16 | 74.18 | 13.72 | 1000.4 | 332.9 | |||||
| 1.75MobileNetV2 | 16/16 | 74.48 | 17.84 | 1326.7 | 441.5 | |||||
| 2.0MobileNetV2 | 16/16 | 74.66 | 22.52 | 1694.5 | 564.1 | |||||
| BASQ [25] | 1.0MobileNetV2 | 2/2 | 64.71 | 0.94 | 17.3 | 6.0 | ||||
| PROM | 0.75MobileNetV2 | (1.58/8)/8 | 64.53 | 1.70 | 11.3 | 3.3 | ||||
| PROM | 1.0MobileNetV2 | (1.58/8)/8 | 68.45 | 1.95 | 15.2 | 4.3 | ||||
| GradScale [41] | 1.0MobileNetV2 | 4/4 | 69.77 | 1.80 | 34.4 | 11.7 | ||||
| PROFIT [34] | 1.0MobileNetV2 | 4/4 | 71.56 | 1.80 | 34.4 | 11.7 | ||||
| BASQ [25] | 1.0MobileNetV2 | 4/4 | 71.98 | 1.80 | 34.4 | 11.7 | ||||
| PROM | 1.25MobileNetV2 | (1.58/8)/8 | 71.58 | 2.60 | 24.5 | 6.2 | ||||
| PROM | 1.5MobileNetV2 | (1.58/8)/8 | 73.17 | 3.31 | 29.5 | 8.1 | ||||
| PROFIT [34] | 1.0MobileNetV2 | 8/8 | 72.35 | 3.54 | 68.7 | 23.1 | ||||
| PROM | 1.5MobileNetV2 | (1.58/8)/8 | 73.17 | 3.31 | 29.5 | 8.1 | ||||
| PROM | 1.75MobileNetV2 | (1.58/8)/8 | 74.24 | 4.10 | 37.5 | 10.1 | ||||
| PROM | 2.0MobileNetV2 | (1.58/8)/8 | 74.81 | 4.96 | 46.4 | 12.4 | ||||
| Method | Model | W/A |
@@
| Baseline | RegNetX-400MF | 16/16 | 72.69 | 10.99 | 650.4 | 215.8 | ||||
| RegNetX-800MF | 16/16 | 75.27 | 14.52 | 1308.8 | 434.7 | |||||
| RegNetX-1.6GF | 16/16 | 76.90 | 18.38 | 2563.9 | 850.8 | |||||
| LSQ [9] | ResNet18 | 2/2 | 67.60 | 2.94 | 97.2 | 34.0 | ||||
| DSQ [13] | ResNet18 | 4/4 | 69.56 | 10.18 | 193.3 | 66.2 | ||||
| BRECQ [27] | ResNet18 | 4/4 | 69.90 | 5.81 | 193.3 | 66.2 | ||||
| PROM | RegNetX-400MF | (1.58/8)/8 | 68.68 | 2.40 | 33.3 | 10.2 | ||||
| PROFIT [34] | MobileNetV1 | 8/8 | 70.16 | 4.25 | 130.5 | 43.7 | ||||
| LSQ [9] | ResNet18 | 8/8 | 71.10 | 11.69 | 386.6 | 131.1 | ||||
| LSQ [9] | ResNet18 | 4/4 | 71.20 | 5.85 | 193.3 | 66.2 | ||||
| LSQ [9] | ResNet34 | 2/2 | 71.60 | 5.48 | 199.6 | 69.7 | ||||
| PROM | RegNetX-800MF | (1.58/8)/8 | 71.96 | 3.01 | 57.7 | 17.3 | ||||
| DSQ [13] | ResNet34 | 4/4 | 72.76 | 15.24 | 396.8 | 135.8 | ||||
| LSQ [9] | ResNet50 | 2/2 | 73.70 | 6.48 | 278.9 | 96.7 | ||||
| PROM | RegNetX-1.6GF | (1.58/8)/8 | 73.91 | 4.15 | 131.1 | 40.3 | ||||
| Method | Model | W/A |
@@
| Baseline | ResNeXt-50(324d) | 16/16 | 77.42 | 50.04 | 7520.0 | 2504.3 | ||||
| ResNeXt-101(328d) | 16/16 | 79.18 | 177.56 | 25567.5 | 8504.9 | |||||
| LSQ [9] | ResNet34 | 8/8 | 74.10 | 21.81 | 793.5 | 269.2 | ||||
| LSQ [9] | ResNet34 | 4/4 | 74.10 | 10.92 | 396.8 | 135.8 | ||||
| BRECQ [27] | ResNet50 | 4/4 | 75.05 | 12.85 | 553.6 | 188.4 | ||||
| LSQ [9] | ResNet50 | 4/4 | 76.70 | 12.85 | 553.6 | 188.4 | ||||
| PROM | ResNeXt-50(324d) | (1.58/8)/8 | 76.71 | 8.97 | 195.1 | 51.0 | ||||
| LSQ [9] | ResNet50 | 8/8 | 76.80 | 25.60 | 1107.3 | 372.6 | ||||
| PROM | ResNeXt-101(328d) | (1.58/8)/8 | 78.21 | 32.09 | 878.8 | 249.6 | ||||
By scaling the quantized model’s width (MobileNetV2) and depth (RegNetX and ResNeXt), we achieve accurate and highly efficient models for varying storage size and energy budgets. For example, we can reduce the energy consumption of a 1.0MobileNetV2 model by a factor of up to 34.4, and the required memory to store it by 3.6. To nearly match the accuracy of the floating-point 1.0MobileNetV2, our 1.25-scaled model requires 2.7 less storage and 23.9 less energy per forward pass, as depicted in Figure 5. Interestingly, our largest model surpasses the floating-point baseline model’s accuracy by 3.03 percentage points, while using 1.4 less storage and 11.9 less energy on architectures.
We observe that our method scales well with the model’s width or depth. Strong quantization in the form of ternary weights for pointwise convolutions heavily restricts the function approximation properties of the model. This issue can be alleviated by providing it with more weights to tune, increasing its representational power. We also note that the 2.0-scaled MobileNetV2 exhibits a much greater amount of ternary weights equal to zero after training when compared to the 1.0 model, as illustrated in Figure 7 of the Supplementary Material. We hypothesize that this indicates an automatic allocation of model capacity during training, essentially "pruning" filters in unimportant layers. As weights which are equal to zero do not need to be computed in the forward pass, the energy expense can be reduced even further on suitable hardware.
| Method | Model | W/A |
[…]
| HAQ [47] | 1.0MobileNetV2 | Mixed | 70.9 | 3.03 | 55.9 | 18.8 | ||||
| Chen et al. [5] | 1.0MobileNetV2 | Mixed | 71.2 | 1.06 | 31.6 | 10.6 | ||||
| FracBits [54] | 1.0MobileNetV2 | Mixed | 71.9 | 2.30 | 28.9 | 9.7 | ||||
| PROM | 1.25MobileNetV2 | (1.58/8)/8 | 71.6 | 2.60 | 24.5 | 6.2 | ||||
| PROM | 1.5MobileNetV2 | (1.58/8)/8 | 73.2 | 3.31 | 29.5 | 8.1 | ||||
We also compare our method to recent mixed-precision approaches [5, 4, 54, 47] that actively learn the bit-width per module in Table 4. The results demonstrate that our principled approach provides a more effective path to model efficiency. For example, our PROM-1.25MobileNetV2 achieves comparable accuracy to the best-performing learned mixed-precision method, while using 36% less energy on a architecture. Moreover, our PROM-1.5MobileNetV2 model surpasses all related methods in accuracy while still maintaining superior energy efficiency. This demonstrates that our fixed, hardware-friendly scheme with ternary weights for costly pointwise convolutions and 8-bit weights for all other layers is a more practical and effective strategy for resource-constrained inference than complex, learned bit-width assignments.
@@ 5.2 Energy Evaluation
As visualized in Figure 1, our method improves the Pareto frontier in the trade-off between top-1 accuracy and inference energy cost for CNN models on ImageNet. In particular, PROM outperforms other methods for quantizing a MobileNetV2 to lower bit-widths.
Our method achieves best-in-category energy efficiency on and architectures while offering similar or better performance than prior work, as demonstrated in Tables 1, 2 and 3. This makes our method well-suited for deployment on battery-powered or resource-constrained devices. Additionally, note that the energy consumption of a model strongly correlates with the ACE metric introduced in [55], which measures a network’s total inference cost on hardware.
@@ 5.4 Comparison to Binary Neural Networks
The concept of low bit-width quantization for convolutional architectures naturally extends to binary networks [37, 22]. Instead of relying on integer addition and multiplication routines, binary networks transform all floating-point operations into logical XNOR and POPCOUNT (counting the number of ones in a bit-string) commands by quantizing both weights and activations into binary tensors.
@@ 5.5 Hardware Support
As quantization is mostly used to decrease the cost of running a model on a given hardware in terms of memory and energy requirements, hardware support for the proposed quantization operations is essential. Most modern CPU and GPU architectures include highly optimized instructions for int8 operations such as addition and multiplication, rendering our proposed method efficient and widely applicable across a variety of hardware.
For models employing 2-bit [13, 9] or 4-bit [13, 9, 27, 34] quantization, an efficient implementation is not straight-forward due to a lack of native hardware support, such as in the popular ARMv8 [2], ARM NEON [1] or modern Intel [23] instruction sets. While specialized hardware could allow these models to run efficiently, general-purpose hardware often lacks optimizations for operations below int8 [52]. This could lead to complicated memory access patterns, as weights need to be packed and unpacked into larger data types, decreasing the model’s efficiency. However, when the hardware support for int4 and int2 operations improves, further reducing the bit-width of the non-ternary components in the network is an interesting direction for future research.
@@ 6 Related Work
Quantizing Weights and Activations to Integers. A common approach to model compression is integer quantization, which converts 32-bit or 16-bit floating-point weights and activations to lower bit-widths. While 8-bit quantization is widely adpoted for its strong performance and hardware support [34, 9, 57], more aggressive 4-bit or 2-bit schemes often suffer from the reduced model capacity and training instabilities [34, 57, 13]. Various techniques have been proposed to mitigate this, including learnable gradient scaling [9], post-training second-order error analysis [27], multi-stage training [13] and selective layer freezing [34]. Another line of work is learned mixed-precision quantization [5, 4, 54, 47], aiming to optimize the bit-widths used in every layer. However, this complicates deployment on general-purpose hardware, as it typically produces a highly fragmented mix of bit-widths. Recently, binary and ternary quantization methods have demonstrated the ability to eliminate expensive multiplications by reducing the forward pass of a module to simple additions [46, 30, 58]. This principle has enabled efficient large language models like BitNet [46] and BitNet b1.58 [30] and has been extended to remove all matrix multiplications from Transformer [45] attention mechanisms [58].
@@
Model Compression via Pruning. Pruning offers another primary path to model compression [51, 18]. Unstructured pruning targets individual low-importance weights, ranging from simple magnitude-based removal [15] to more sophisticated methods for identifying redundancy [50, 49]. For practical speedups on general-purpose hardware, structured pruning is used to remove entire filters or channels. These methods can be static, resulting in a fixed architecture [48, 26], or dynamic, using gates to toggle components based on input complexity [12, 8]. Recent work has even demonstrated instant, training-free compression [33]. While pruning reduces parameter counts and can be paired with quantization for greater efficiency [15], exploring this synergy is beyond the scope of this work, but remains a compelling avenue for future research.